OpenStudy (anonymous):

In Kinematic Equations, I understand the variables, I can even understand them when they are constructed in a couple ways, namely, Vf=at + Vi and d=Vi*t + 1/2a*t2; however, I do not understand what the variables "mean" in the other equations. Can anyone help?

OpenStudy (anonymous):

yes Vf is final velocity and Vi is initial velocity. a = acceleration, d = distance and t= time

OpenStudy (anonymous):

what 'other' variables do you mean?

OpenStudy (anonymous):

Variables they are name using the same system but u could name them wherever you want : example of those convention naming variable is the use of S or x for the same variable all depend on the system use. u could use v=s/t or v=x/t or v=y/t . x mean u are using and axis xy to name those variable on the axis . like in algebra u could name variable either way u want But always keep the same system.